OpenStudy (anonymous):

simplify: 10^4(10^3÷10^2)

OpenStudy (anonymous):

\[10^4.{10^3 \over 10^2}=10^4(10)=10^5\]

OpenStudy (anonymous):

The idea here is when you multiply two numbers that have the same base, you add the exponents, and when you divide them you, you subtract exponents. Does that make sense?

OpenStudy (anonymous):

simplify: (-2x^2)(-3xy)(-y)

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

10^5 is ans

OpenStudy (anonymous):

-6x^3y^2